OpenStudy (anonymous):

i need help the question is Explain wether or not the expression x^2 +24x +144 fits one of the special patterns if so which one? i have no clue

OpenStudy (anonymous):

like the difference of two squares

OpenStudy (anonymous):

Well not "difference of squares" but it sure does fit the "special case". What that is is like, if you have something in the form of \(ax^2+bx+c\), and \(a\) and \(c\) are both squares, then you might be able to write the whole thing as \((\sqrt{a}(x)+\sqrt{c})^2\) So in this case, our \(a=1=1^2\) and \(c=144=12^2\). So we can write the whole thing as \((\sqrt{1}(x)+\sqrt{144})^2=(1(x)+12)^2=(x+12)^2\)